# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

### v2

@@ A.6 Upper bound on the second moment of the ratio
$m_{1}(\Phi_{t})$ $\leq e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}m_{1}(\Phi_{0})$
$~{}~{}~{}+2\int_{0}^{t}\left[\|\nabla\log(p_{s}^{V})(0)\|m_{1}(\Phi_{s})+\sqrt{dL}(m_{0}(\Phi_{s})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\right]\frac{e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}}{e^{s(d+1)}(L(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s\,.$
Using Lemma 10, we have that $m_{0}(\Phi_{s})\leq\frac{Z_{V}^{2}}{Z_{2V}}e^{sd}(L(1-e^{-2s})+e^{-2s})^{d}$ hence the first term of the integral is upper-bounded as:
$\int_{0}^{t}\frac{e^{-s(d+1)}m_{0}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|}{(L(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s\leq\frac{Z_{V}^{2}}{Z_{2V}}\int_{0}^{t}\|\nabla\log(p_{s}^{V})(0)\|\frac{e^{-s}}{L(1-e^{-2s})+e^{-2s}}\mathrm{d}s\,.$
@@
$\int_{0}^{t}\frac{\|\nabla\log(p_{s}^{V})(0)\|e^{-s}}{L(1-e^{-2s})+e^{-2s}}\mathrm{d}s\leq\sqrt{\int_{0}^{t}\|\nabla\log(p_{s}^{V})(0)\|^{2}\mathrm{d}s}\sqrt{\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s}\,.$
@@
$\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s$ $=\frac{1}{2}\int_{e^{-2t}}^{1}\frac{1}{(L(1-u)+u)^{2}}\mathrm{d}u$
$=\frac{1}{2(1-L)}[-\frac{1}{L(1-u)+u}]_{e^{-2t}}^{1}$
$=\frac{1}{2(1-L)}(\frac{1}{L(1-e^{-2t})+e^{-2t}}-1)$
$=\frac{1-e^{-2t}}{2}$
@@
$\int_{0}^{t}m_{0}(\Phi_{s})^{\frac{d-1}{d}}\frac{e^{-s(d+1)}}{(L(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s$ $\leq\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s$
$=\frac{1-e^{-2t}}{2}$
@@
$m_{1}(\Phi_{t})\leq\frac{Z_{V}^{2}}{Z_{2V}}e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}\left(\frac{Z_{2V}}{Z_{V}^{2}}m_{1}(\Phi_{0})+\sqrt{-2\log(\mu(0))+d\log\left(\frac{L}{2\pi}\right)}+\sqrt{dL}\right)\,.$
@@
We can derive our upper-bound on $m_{2}(\Phi_{t})$.
@@
$\partial m_{2}(\Phi_{t})$ $=\int\|z\|^{2}\partial\Phi_{t}\mathrm{d}z$
@@
$=2dm_{0}(\Phi_{t})+dm_{2}(\Phi_{t})+2m_{2}(\Phi_{t})+2\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z$
$~{}~{}+4\int\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),z\rangle\Phi_{t}\mathrm{d}z\,.$
The first term $\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z$ is upper-bounded by $\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}dm_{2}(\Phi_{t})$ and for the second term we have
@@
$\leq\int\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}\|z\|^{2}\Phi_{t}\mathrm{d}z+\|\log(p_{t}^{V})(0)\|\int\|z\|\Phi_{t}\mathrm{d}z$
$=\frac{(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}m_{2}(\Phi_{t})+\|\nabla\log(p_{t}^{V})(0)\|m_{1}(\Phi_{t})\,.$
@@
$\partial m_{2}(\Phi_{t})\leq(d+2)\left(1+\frac{2(L-1)e^{-2t}}{(1-e^{-2t})(L-1)+1}\right)m_{2}(\Phi_{t})+4\|\nabla\log(p_{t}^{V})(0)\|m_{1}(\Phi_{t})+2dm_{0}(\Phi_{t})\,.$
@@
$m_{2}(\Phi_{t})\leq e^{t(d+2)}(L(1-e^{-2t})+e^{-2t})^{d+2}\left(m_{2}(\Phi_{0})+2\int_{0}^{t}\frac{e^{-s(d+2)}(2m_{1}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|+dm_{0}(\Phi_{s}))}{(L(1-e^{-2s})+e^{-2s})^{(d+2)}}\mathrm{d}s\right)$
Recalling the upper-bound $m_{1}(\Phi_{t})\leq e^{t(d+1)}(L(1-e^{-2t})+e^{-2t})^{d+1}C$ where $C$ is defined in Lemma 24, we upper-bound the integral term as
$\int_{0}^{t}\frac{e^{-s(d+2)}m_{1}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|}{(L(1-e^{-2s})+e^{-2s})^{(d+2)}}\mathrm{d}s$ $\leq C\int_{0}^{t}\frac{e^{-s}}{L(1-e^{-2s})+e^{-2s})}\|\nabla\log(p_{s}^{V})(0)\|\mathrm{d}s$
$\leq C\sqrt{\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s}\sqrt{\int_{0}^{t}\|\nabla\log(p_{t}^{V})(0)\|^{2}\mathrm{d}s}$
@@
$\int_{0}^{t}\frac{e^{-s(d+2)}}{(L(1-e^{-2s})+e^{-2s})^{(d+2)}}m_{0}(\Phi_{s})\mathrm{d}s$ $\leq\int_{0}^{t}\frac{e^{-2s}}{(L(1-e^{-2s})+e^{-2s})^{2}}$
@@
$m_{2}(\Phi_{t})\leq e^{t(d+2)}(L(1-e^{-2t})+e^{-2t})^{d+2}\left(m_{2}(\Phi_{0})+2C\sqrt{-2\log\left(\mu(0)\right)+d\log(\frac{L}{2\pi})}+d\right)\,.$
@@
$m_{2}(\Phi_{t})\leq\frac{2e^{t(d+2)}Z_{V}^{2}}{Z_{2V}}(L(1-e^{-2t})+e^{-2t})^{d+2}\left[m_{2}(\mu)+d(L+1)-4\log\left(\mu(0)\right)+2d\log(\frac{L}{2\pi})\right]\,.$
A.7 Integrated average squared error of the estimator
Proposition 25
Under Assumptions 1-2 and the additional assumptions that $\|\nabla\log(p_{t}^{V})(0)\|,\|\nabla\log(p_{t}^{2V})(0)\|\leq\xi$ for some $\xi\geq 0$ and that $\mu$ has finite second moment $m_{2}(\mu)$, it holds that
$\begin{split}\int\mathbb{E}[\|\hat{s}_{n,t}(z)-\nabla\log(p_{t})(z)\|^{2}]\mathrm{d}p_{t}^{V}(z)\lesssim~{}&\frac{e^{2t(d+1)}(d+\log(n))(L+1)^{d}}{n(1-e^{-2t})^{2}}(K_{1}+K_{2})\,,\end{split}$
where we defined
$\begin{cases}&K_{1}=(L+\beta+1)^{2}(L+1)^{2}\bigg{[}m_{2}(\mu)+d(L+1)+2d\log\left(\frac{L\mu(0)^{-2/d}}{2\pi}\right)\bigg{]}\,,\\ &K_{2}=d\beta+\xi^{2}+1\,.\end{cases}$ (10)

### v3

@@ Appendix E Proof of Lemma 7
$m_{1}$ $(\Phi_{t})\leq e^{t(d+1)}(\beta(1-e^{-2t})+e^{-2t})^{d+1}m_{1}(\Phi_{0})$
$+2\int_{0}^{t}\left[\|\nabla\log(p_{s}^{V})(0)\|m_{1}(\Phi_{s})+\sqrt{d\beta}(m_{0}(\Phi_{s})Z_{2V}/Z_{V}^{2})^{\frac{d-1}{d}}\right]\frac{e^{t(d+1)}(\beta(1-e^{-2t})+e^{-2t})^{d+1}}{e^{s(d+1)}(\beta(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s\,.$
Using Lemma 7, we have that $m_{0}(\Phi_{s})\leq\frac{Z_{V}^{2}}{Z_{2V}}e^{sd}(\beta(1-e^{-2s})+e^{-2s})^{d}$ hence the first term of the integral is upper-bounded as:
$\int_{0}^{t}\frac{e^{-s(d+1)}m_{0}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|}{(\beta(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s\leq\frac{Z_{V}^{2}}{Z_{2V}}\int_{0}^{t}\|\nabla\log(p_{s}^{V})(0)\|\frac{e^{-s}}{\beta(1-e^{-2s})+e^{-2s}}\mathrm{d}s\,.$
@@
$\int_{0}^{t}\frac{\|\nabla\log(p_{s}^{V})(0)\|e^{-s}}{L(1-e^{-2s})+e^{-2s}}\mathrm{d}s\leq\sqrt{\int_{0}^{t}\|\nabla\log(p_{s}^{V})(0)\|^{2}\mathrm{d}s}\sqrt{\int_{0}^{t}\frac{e^{-2s}}{(\beta(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s}\,.$
@@
$\int_{0}^{t}\frac{e^{-2s}}{(\beta(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s$ $=\frac{1}{2}\int_{e^{-2t}}^{1}\frac{1}{(\beta(1-u)+u)^{2}}\mathrm{d}u$
$=\frac{1}{2(1-\beta)}[-\frac{1}{\beta(1-u)+u}]_{e^{-2t}}^{1}$
$=\frac{1}{2(1-\beta)}(\frac{1}{\beta(1-e^{-2t})+e^{-2t}}-1)$
$=\frac{1-e^{-2t}}{2(\beta(1-e^{-2t})+e^{-2t})}$
@@
$\int_{0}^{t}m_{0}(\Phi_{s})^{\frac{d-1}{d}}\frac{e^{-s(d+1)}}{(\beta(1-e^{-2s})+e^{-2s})^{d+1}}\mathrm{d}s$ $\leq\int_{0}^{t}\frac{e^{-2s}}{(\beta(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s$
$=\frac{1-e^{-2t}}{2(\beta(1-e^{-2t})+e^{-2t})}$
@@
$m_{1}(\Phi_{t})\leq\frac{Z_{V}^{2}}{Z_{2V}}e^{t(d+1)}(\beta(1-e^{-2t})+e^{-2t})^{d+1}\left(\frac{Z_{2V}}{Z_{V}^{2}}m_{1}(\Phi_{0})+\sqrt{-2\log(\mu(0))+d\log\left(\frac{\beta}{2\pi}\right)}+\sqrt{d\beta}\right)\,.$
@@
We can now derive our upper-bound on $m_{2}(\Phi_{t})$.
@@
$\partial_{t}m_{2}(\Phi_{t})$ $=\int\|z\|^{2}\partial_{t}\Phi_{t}\mathrm{d}z$
@@
$=2dm_{0}(\Phi_{t})+dm_{2}(\Phi_{t})+2m_{2}(\Phi_{t})$
$\qquad+2\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z+4\int\langle\nabla\log(\pi)-\nabla\log(p_{t}^{V}),z\rangle\Phi_{t}\mathrm{d}z\,.$
The first term $\int(\Delta\log(\pi)-\Delta\log(p_{t}^{V}))\|z\|^{2}\Phi_{t}\mathrm{d}z$ is upper-bounded by $\frac{(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}dm_{2}(\Phi_{t})$ and for the second term we have
@@
$\leq\int\frac{(L-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}\|z\|^{2}\Phi_{t}\mathrm{d}z+\|\log(p_{t}^{V})(0)\|\int\|z\|\Phi_{t}\mathrm{d}z$
$=\frac{(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}m_{2}(\Phi_{t})+\|\nabla\log(p_{t}^{V})(0)\|m_{1}(\Phi_{t})\,.$
@@
$\partial_{t}m_{2}(\Phi_{t})\leq(d+2)\left(1+\frac{2(\beta-1)e^{-2t}}{(1-e^{-2t})(\beta-1)+1}\right)m_{2}(\Phi_{t})+4\|\nabla\log(p_{t}^{V})(0)\|m_{1}(\Phi_{t})+2dm_{0}(\Phi_{t})\,.$
@@
$m_{2}(\Phi_{t})\leq e^{t(d+2)}(\beta(1-e^{-2t})+e^{-2t})^{d+2}\left(m_{2}(\Phi_{0})+2\int_{0}^{t}\frac{e^{-s(d+2)}(2m_{1}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|+dm_{0}(\Phi_{s}))}{(\beta(1-e^{-2s})+e^{-2s})^{(d+2)}}\mathrm{d}s\right)$
Recalling the upper-bound $m_{1}(\Phi_{t})\leq e^{t(d+1)}(\beta(1-e^{-2t})+e^{-2t})^{d+1}C$ where $C$ is defined in Lemma 19, we upper-bound the integral term as
$\int_{0}^{t}\frac{e^{-s(d+2)}m_{1}(\Phi_{s})\|\nabla\log(p_{s}^{V})(0)\|}{(\beta(1-e^{-2s})+e^{-2s})^{(d+2)}}\mathrm{d}s$ $\leq C\int_{0}^{t}\frac{e^{-s}}{\beta(1-e^{-2s})+e^{-2s})}\|\nabla\log(p_{s}^{V})(0)\|\mathrm{d}s$
$\leq C\sqrt{\int_{0}^{t}\frac{e^{-2s}}{(\beta(1-e^{-2s})+e^{-2s})^{2}}\mathrm{d}s}\sqrt{\int_{0}^{t}\|\nabla\log(p_{t}^{V})(0)\|^{2}\mathrm{d}s}$
@@
$\int_{0}^{t}\frac{e^{-s(d+2)}}{(\beta(1-e^{-2s})+e^{-2s})^{(d+2)}}m_{0}(\Phi_{s})\mathrm{d}s$ $\leq\int_{0}^{t}\frac{e^{-2s}}{(\beta(1-e^{-2s})+e^{-2s})^{2}}$
@@
$m_{2}(\Phi_{t})\leq e^{t(d+2)}(\beta(1-e^{-2t})+e^{-2t})^{d+2}\left(m_{2}(\Phi_{0})+2C\sqrt{-2\log\left(\mu(0)\right)+d\log(\frac{\beta}{2\pi})}+d\right)\,.$
@@
$m_{2}(\Phi_{t})\leq\frac{2e^{t(d+2)}Z_{V}^{2}}{Z_{2V}}(\beta(1-e^{-2t})+e^{-2t})^{d+2}\left[m_{2}(\mu)+d(\beta+1)-4\log\left(\mu(0)\right)+2d\log(\frac{\beta}{2\pi})\right]\,.$
Appendix F Proof of Theorem 9